A Complete Bibliography of Publications in BJHS Themes

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.07

Title word cross-reference

1830s [Som18]. 1970s [Pet17].

absence [Som21]. Accepting [Gho16]. across [LS16]. African [Lan17]. after [DM17]. afterlife [Kow19]. age [Rad17]. agencies [Ree17b]. agency [Jon17, Nel17, Rad17, Sah17]. agents [Ree17a]. agricultural [Say17]. alchemical [Mar20b]. alchemically [Ram20]. Alfred [Hop19]. already [GMPW16]. Alto [GMPW16]. America [Kak19, dO18]. anatomical [Hop19]. Ancient [Mar20b, Kak19, Kak19]. Animal [CDSW17, Rad17, Ree17a, HB17, Jon17, Nel17, Ree17b, Sah17]. antagonisms [Zak21]. Anthropocene [Ree17a]. anthropological [Roq19]. appropriation [dO18]. arrives [GMPW16]. ascent [Jia21]. Asia [GMPW16]. Asian [Muk21]. Atlantic [DM17]. automation [Dic20].

Babel [Kap21]. Back [Ano16a, Ano17a, Ano18a, Ano19a, Ano20a, Ano21a]. bank [Cur19]. bawdy [Ham21]. bean [Cur19]. becomings [HB17]. Bengali [Roy18]. Benninghoff [Hop19]. beyond [Mei20]. bird [Muk21]. BJT [Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a]. blood [Roq19]. book [CGL20, Han20, Jar20, Leo20, Mar20a, MWP20]. border [LS16]. Bovine [HB17]. boy [Bow18]. Britain [Bow18]. British [FM16, Sah17, Say17]. built [Mei20]. bulbuls [Muk21]. Burma [Sah17]. c [CDSW17]. calculus [Pet17]. Cambridge [Jar19]. canon [Che20]. capital [Sah17]. capitals [Wu16]. card [Roq19]. care [Ban19a, Ban19b]. case [Wu16]. cat [Pet17]. Catalogues [Por19]. categories [Zak21]. centres [Ban19a, Ban19b]. centuries [Che20, Han20]. century [Bow18, Gro20, PL16, Ree17b, Say17, Zil18, dO18]. Chapters [Che20]. Charles [Hal21, RS21]. chemical [Gro20]. Children [dO18, Roy18, ZN18]. chimpanzees [Lan17]. China [FM16, Jia21, LS16, PW16, PL16, Roc16, SS16, Sid16]. Chinese [DL16, Han20, Wu16]. City [Pet17]. classifying [Che20]. Cloning [Cre20]. coal [Wu16]. Coded [Dic20]. Cold [Nie18]. collages [Lóp19]. collect [Del19]. collection [Cur19, Por19]. collections [Ban19a, Ban19b, Hop19, JKB19, Lóp19, Mar20b, Roq19]. colonial [Roq19, Roy18]. Colonizing [Sah17]. commentaries [Che20]. Commentary [Del19, Rea19]. common [Gho16]. conduct [Dic20]. conservation [Cur19, Jon17, LS16]. Conserving [DM17]. Constitutions [Set21]. context [Roy18]. contexts [Wu16]. conversation [GMPW16]. Count [Hop19]. country [Zil18]. countryside [Zil18]. Cover [Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a]. craft [Jar20]. crosses [MWP20]. cultures [Muk21]. curation [Ban19a, Ban19b].

dairy [HB17]. Darwin [Hal21, Ham21, Jia21, MS21, Muk21, RS21, Set21, Som21]. Darwinian [Mil21]. Database [SW19]. decolonial [Kow19]. deep [Roc16]. deletion [SW19]. Denny [RS21]. Descent [MS21, Som21, Ham21, Jia21]. destruction [SW19]. did [Nel17, Som21]. difference [Gho16]. different [Wu16]. disciplinary [Por19]. discursive [Wu16]. DNA [SW19, Cre20]. do [Lan17]. dogs [Nel17]. double [Kow19]. Drosophila [Ban19a, Ban19b]. during [Say17].

earlier [GMPW16]. early

[Bow18, Jar20, Leo20, Mar20a, Ram20, Zil18]. educational [Gom18]. Egyptian [Mar20b]. elbow [Han20]. elephants [Sah17]. emeritus [Hop19]. Encyclopedia [dO18]. Encyclopedic [Gro20]. end [JKB19, Kak19]. ending [SW19]. ends [Rea19]. engagement [Roc16]. England [Ram20]. entropic [Por19]. environmental [Jon17]. epistemologies [SS16]. Ernst [Mei20]. Erratum [Ban19b]. essay [Phi16]. ethics [Hal21]. everlasting [Rea19]. evidence [RS21]. evolution [Hal21, Jia21, Mil21]. Evolutionary [Zak21]. example [Rad17]. exchanges [Gho16]. exhaustion [Por19]. experimental [Jar19]. experimentation [Gom18]. exploring [Jia21].

farm [Say17]. fates [Hop19]. Ferdinand [Hop19]. fiction [GMPW16]. fictionalizing [Zil18]. field [Lan17, Ree17b, Say17]. fieldwork [Mar20a]. fierce [Jon17]. fighting [Muk21]. fire [Jon17]. flight [Nel17]. formatting [Mei20]. Foster [Kap21]. fourteenth [Han20]. fourth [Han20]. Front [Ano16b, Ano17b, Ano18b, Ano19b, Ano20b, Ano21b]. future [GMPW16].

gains [Por19]. Gander [DM17].

Geiseltalmuseum [Lóp19]. gendered [Ham21]. genre [Som18]. giants [PL16]. Glasgow [Por19]. global [Nie18, Sid16]. Graeco [Mar20b]. great [Pet17]. Green [SS16, Jon17]. Green-revolution [SS16]. ground [Gho16]. guides [Ram20].

H [Rad17]. handbooks

[CGL20, Gro20, Mar20b]. heirloom [Cur19]. Henry [RS21]. High [DL16, GMPW16].

High-tech [DL16]. historical [Jar19]. Histories [Phi16, GMPW16, HB17]. historiography [Roc16, Ree17a]. history [CDSW17, CGL20, Cre20, FM16, Gom18, Ree17a, Ree17b, Sid16, Som18]. human [HB17, Jia21, MS21, RS21]. humans [Lan17]. Hunterian [Por19]. hybrid [MWP20].

Ideas [Nie18]. identity [SS16]. ideological [Wu16]. imperial [Sah17]. India [FM16, LS16, PW16, PL16, Roc16, SS16, Sid16]. Indian [DL16, Gho16, Kak19, Wu16]. ing [Jon17]. institution [Mar20a]. instructions [Che20]. instrument [Jar20]. integrated [Nie18]. intentionality [Ree17b]. Investigating [Wu16].

Japan [Mar20a]. Japanese [Lan17]. Joseph [Roc16]. juvenile [Roy18].

Kennewick [Kak19]. knowledge [Gro20, Mar20a, SS16].

lab [Jar19]. laboratories [JKB19]. Laboratory [Cre20, Lan17]. language [Kap21]. late [Jia21]. Latin [dO18]. laws [Gom18]. learned [MWP20]. Learning [CGL20, Leo20, Han20]. Lebensgestaltungslehre [Mei20]. LeCron [Kap21]. leopard [LS16]. liberal [DL16]. lice [RS21]. life [Gro20, Mei20, Som18]. Living [Ban19a, Ban19b]. London [Leo20]. loss [JKB19]. losses [Por19]. love [Roc16].

MACSYMA [Dic20]. Magazine [Bow18]. magazines [Roy18]. make [Som21]. making [Dic20, She21, Sid16]. Man [Kak19, Ham21, Som21]. Man/the [Kak19]. management [Say17]. manned [Nel17]. Manual [Cre20]. manuals [CGL20, Leo20]. Mary [Kap21]. Mathematical [Che20]. mathematics [Dic20, Jar20]. matter [Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a]. meaning [JKB19, Som21]. Meccano [Bow18]. medicine [CDSW17, Han20, Leo20, Set21]. memoir [Hop19]. memory [Jon17]. men [RS21]. Mendelians [MWP20]. metaphors [Han20]. milking [HB17]. missions [Roq19]. modern [Jar20, Leo20, Mar20a, Ram20, Say17, Rad17]. Molecular [Cre20]. movements [Pet17]. Museum [Por19, Jar19, Kow19]. museums [JKB19]. mutilation [Pet17].

National [Jon17, SW19]. natural [FM16, Gom18]. natural-history [Gom18]. nature [DM17, Mar20a, MS21, Wu16]. Needham [Roc16]. Negotiating [FM16]. neo [DL16]. neo-liberal [DL16]. Neufert [Mei20]. Newfoundland [DM17]. Nine [Che20]. non [Kak19]. non-Indian [Kak19]. noon [GMPW16]. North [Kak19]. novels [Zil18].

objects [JKB19]. Observation [Gom18]. One [Kak19]. origins [Kap21]. other [Hal21]. other-regarding [Hal21].

Palaeoamericans [Kak19]. Palaeontological [Som18]. palm [Han20]. Palo [GMPW16]. Park [Jon17]. parks [DL16]. past [Che20]. peasant [SS16]. pedagogues [Som18]. people [Zil18, ZN18]. philosophical [Ram20]. Photo [Phi16]. Planning [PW16]. pointing [Han20]. Police [SW19]. politics [Nie18]. popular [Ham21]. popularization [Bow18]. Portugal [Zil18]. Portuguese [Gom18]. postcolonial [Kow19]. practical [Jar20]. practice [Jar19, Ram20, Ree17b]. practices [Nie18]. prehistory [Som18]. primate [Som21]. primatology [Lan17]. Procedures [Che20]. production [SS16]. programme [Nel17]. progress [Zak21]. prop [Kow19]. Punnett [MWP20]. race [MS21, Set21, She21]. races [RS21]. radical [Ham21]. rats [Say17]. Reading [Che20, Ram20, Leo20]. reception [Ham21]. Recipes [Cre20, Mar20b]. reclassifying [Che20]. recombining [Cre20]. reconceptualizing [LS16]. records [SW19]. regarding [Hal21]. reimagining [GMPW16]. remains [Roq19]. remaking [HB17]. repatriation [Kak19]. Restor [Jon17]. revolution [SS16]. river [DM17]. robotic [HB17]. role [Jia21]. roles [CDSW17].

salmon [DM17]. samples [SW19]. schools [Gom18]. Science [PL16, Roy18, dO18, Bow18, CGL20, DL16, GMPW16, Nie18, PW16, Ree17a, Sah17, Say17, Sid16, Zil18, ZN18]. sciences [Gro20, Jar19]. scientific [SS16]. secondary [Gom18]. seed [Cur19]. seeking [Gho16]. selection [Hal21, Muk21, Set21, She21]. sex [MS21, Pet17]. Sexual [She21, Hal21, Muk21]. sexualities [Mil21]. Sino [Gho16]. snow [LS16]. social [Pet17]. South [Muk21]. Soviet [Nel17]. space [Nel17, Sid16]. species [RS21]. Speculative [Phi16]. Spee [Hop19]. Spencer [Kow19]. squares [MWP20]. statistical [Gho16]. stock [Ban19a, Ban19b]. studies [Ree17b, Wu16]. Studying [LS16]. subjectivity [HB17]. surgical [Leo20]. Synthesis [Rad17]. Synthetic [Lan17]. systems [HB17].

Taking [Zil18]. taphonomy [Lóp19]. teaching [Gom18, Nie18]. tech [DL16]. technique [Jar20]. technocracy [SS16]. technologies [HB17]. technology [GMPW16, PW16]. textual [Mar20a]. their [MWP20]. them [Che20]. theory [Muk21, Ree17a]. there [GMPW16]. third [Che20]. thirteenth [Che20]. Thorpe [Rad17]. three [Zak21]. tomorrow [GMPW16]. Total [Gro20]. toys [Bow18]. traces [CDSW17]. trade [MWP20]. tragedy [Hop19]. traits [Zak21]. transformations [Cur19]. transitional [FM16]. tree [Som21]. turn [DL16]. twentieth [Bow18, Gro20, PL16, Ree17b, Say17, Zil18, dO18]. twentieth-century [Bow18, Gro20, Ree17b, Zil18].

UK [SW19]. undead [Sah17]. Unravelling [Kap21]. users [Dic20]. utilitarian [Pet17]. utopianism [DL16].

varieties [RS21]. vegetable [Cur19].

versus [Gom18]. volume [Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a].

W. [Rad17]. War [Nie18]. Was [SW19]. Wildlife [Ree17b]. within [Wu16]. wolf [Jon17]. women [Jia21]. Worlds [ZN18]. writing [GMPW16, Leo20].

Yellowstone [Jon17]. York [Pet17]. young [Zil18, ZN18].

References

Anonymous:2016:BVCb

[Ano16a] Anonymous. BJT volume 1 cover and back matter. BJHS Themes, 1:b1–b8, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                                  URL https://www.cambridge.
                                                  org/core/product/
                                                  8C15E2C2AE512C25836CAEF78EC05BA0.

Anonymous:2016:BVCa

[Ano16b] Anonymous. BJT volume 1 cover and front matter. BJHS Themes, 1:f1–f2, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
C15494106809B35EDEB7A697926F2833.

Anonymous:2017:BVCb

[Ano17a] Anonymous. BJT volume 2 cover and back matter. BJHS Themes, 2:b1–b2, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
129297FF22E1F174432E9DA0B2F76798.

Anonymous:2017:BVCa

[Ano17b] Anonymous. BJT volume 2 cover and front matter. BJHS Themes, 2:f1–f2, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
FA72CC5794567652B484E4113A638221.

Anonymous:2018:BVCb

[Ano18a] Anonymous. BJT volume 3 cover and back matter. BJHS Themes, 3:b1–b3, 2018. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/bjt-volume-
3-cover-and-back-matter/
AE2124148C9E741061C1C659E640C1FC.

Anonymous:2018:BVCa

[Ano18b] Anonymous. BJT volume 3 cover and front matter. BJHS Themes, 3:f1–f2, 2018. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-

                                        themes/article/bjt-volume-
                                        3-cover-and-front-matter/
                                        F7F58B243D78E9CFCC687C761CE3ECE0.

Anonymous:2019:BVCb

[Ano19a] Anonymous. BJT volume 4 cover and back matter. BJHS Themes, 4:b1–b2, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/bjt-volume-
                                        4-cover-and-back-matter/
                                        48EA9206F47DB1621E5DAEE48DD26F57.

Anonymous:2019:BVCa

[Ano19b] Anonymous. BJT volume 4 cover and front matter. BJHS Themes, 4:f1–f2, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/bjt-volume-
                                        4-cover-and-front-matter/
                                        EC854EBF2CBC03F5D8EF885BE8627EAE.

Anonymous:2020:BVCb

[Ano20a] Anonymous. BJT volume 5 cover and back matter. BJHS Themes, 5::b1–b2, 2020. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/bjt-volume-
                                        5-cover-and-back-matter/
                                        3D0CA317082021622922DFE712904631.

Anonymous:2020:BVCa

[Ano20b] Anonymous. BJT volume 5 cover and front matter. BJHS Themes, 5::f1–f2, 2020. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/bjt-volume-
5-cover-and-front-matter/
3DC0332EDED3A74150485A3C6FAFA7A8.

Anonymous:2021:BVCb

[Ano21a] Anonymous. BJT volume 6 cover and back matter. BJHS Themes, 6:b1–b2, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/bjt-volume-
6-cover-and-back-matter/
5DE7C5E0B3615EC65A480F38A6BD98BE.

Anonymous:2021:BVCa

[Ano21b] Anonymous. BJT volume 6 cover and front matter. BJHS Themes, 6:f1–f2, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/bjt-volume-
6-cover-and-front-matter/
365EE900A5FE5959D708A80D226B51B9.

Bangham:2019:LCCa

[Ban19a] Jenny Bangham. Living collections: care and curation at Drosophila stock centres. BJHS Themes, 4:123–147, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/living-
collections-care-and-

                                        curation-at-drosophila-
                                        stock-centres/
                                        08E50B05EB5EF869DFB969D7A77E611E.

Bangham:2019:LCCb

[Ban19b] Jenny Bangham. Living collections: care and curation at Drosophila stock centres — Erratum. BJHS Themes, 4:293, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/living-
                                        collections-care-and-
                                        curation-at-drosophila-
                                        stock-centres-erratum/
                                        C9F1F27F023B53F0CC6D730295E67825.

See erratum [Ban19b].

Bowler:2018:MMB

[Bow18] Peter J. Bowler. Meccano Magazine: boys’ toys and the popularization of science in early twentieth-century Britain. BJHS Themes, 3:129–146, 2018. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/meccano-
                                        magazine-boys-toys-and-the-
                                        popularization-of-science-
                                        in-early-twentiethcentury-
                                        britain/
                                        2A15FBD05B6FB697BFC1C12456B10FB9.

Cassidy:2017:ART

[CDSW17] Angela Cassidy, Rachel Mason Dentinger, Kathryn Schoefert, and Abigail Woods. Animal roles and traces in the history of medicine, c. 1880–1980. BJHS Themes, 2:11–33, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
C988416768BA334B73A708F2B4189F4E.

Creager:2020:LBM

[CGL20] Angela N. H. Creager, Mathias Grote, and Elaine Leong. Learning by the book: manuals and handbooks in the history of science. BJHS Themes, 5::1–13, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/learning-by-the-
book-manuals-and-handbooks-
in-the-history-of-science/
C79B348D396B85057826406507192D51.

Chemla:2020:RIP

[Che20] Karine Chemla. Reading instructions of the past, classifying them, and reclassifying them: commentaries on the canon The Nine Chapters on Mathematical Procedures from the third to the thirteenth centuries. BJHS Themes, 5::15–37, 2020. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/reading-
instructions-of-the-past-
classifying-them-and-
reclassifying-them-
commentaries-on-the-canon-
the-nine-chapters-on-
mathematical-procedures-
from-the-third-to-the-

                                        thirteenth-centuries/
                                        F402B3DE252D1799135618C259E5C706.

Creager:2020:RRD

[Cre20] Angela N. H. Creager. Recipes for recombining DNA: A history of Molecular Cloning: A Laboratory Manual. BJHS Themes, 5::225–243, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/recipes-for-
                                        recombining-dna-a-history-
                                        of-molecular-cloning-a-
                                        laboratory-manual/
                                        F4EE6A7FFA4991B714A33D474BC1CF76.

Curry:2019:BCS

[Cur19] Helen Anne Curry. From bean collection to seed bank: transformations in heirloom vegetable conservation, 1970–1985. BJHS Themes, 4:149–167, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/from-bean-
                                        collection-to-seed-bank-
                                        transformations-in-
                                        heirloom-vegetable-
                                        conservation-19701985/
                                        1D973B01429B782FDB86B9206559C456.

Delbourgo:2019:CC

[Del19] James Delbourgo. Commentary: collect or die. BJHS Themes, 4:273–281, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-

themes/article/commentary-
collect-or-die/
565687DA16528431AF0D355C876E6F6A.

Dick:2020:CCM

[Dic20] Stephanie A. Dick. Coded conduct: making MACSYMA users and the automation of mathematics. BJHS Themes, 5::205–224, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/bjhs-themes/
article/coded-conduct-
making-macsyma-users-and-
the-automation-of-
mathematics/
19936AF1E826C13BA410EFDB704026DF.

Das:2016:HTU

[DL16] Diganta Das and Tong Lam.

High-tech utopianism: Chinese and Indian science parks in the neo-liberal turn. BJHS Themes, 1:221–238, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
541DB583136CDA6BAE74FE1E2B8B91F2.

Daniels:2017:CAS

[DM17] Jennifer Daniels and Charles Mather. Conserving Atlantic salmon ‘after nature’ on Newfoundland’s Gander river. BJHS Themes, 2:191–213, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
01D76FCF25D98BD9C86FC2216924C6FA.

deOliveira:2018:SCE

[dO18] Bernardo Jefferson de Oliveira. Science in The Children’s Encyclopedia and its appropriation in the twentieth century in Latin America. BJHS Themes, 3: 105–128, 2018. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/science-in-the-
                                        childrens-encyclopedia-and-
                                        its-appropriation-in-the-
                                        twentieth-century-in-latin-
                                        america/
                                        DEBF2855F1B136D02324EA98CB1596C3.

Fan:2016:NNH

[FM16] Fa-Ti Fan and John Mathew. Negotiating natural history in transitional China and British India. BJHS Themes, 1:43–59, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        DE38EAA9231225BBC09B8230F02752FC.

Ghosh:2016:ADS

[Gho16] Arunabh Ghosh. Accepting difference, seeking common ground: Sino–Indian statistical exchanges 1951–1959. BJHS Themes, 1:61– 82, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        823FAB90F5CCCB156A384EB0D3A55EAE.

Greenspan:2016:FAE

[GMPW16] Anna Greenspan, Anil Menon, Kavita Philip, and Jeffrey Wasserstrom. The future arrives earlier in Palo Alto (but when it’s high noon there, it’s already tomorrow in Asia): a conversation about writing science fiction and reimagining histories of science and technology. BJHS Themes, 1:249–266, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
2161E640FDBAB95BCECC52C9BBA0C3C7.

Gomes:2018:OVE

[Gom18] Inês Gomes. Observation versus experimentation in naturalhistory teaching in Portuguese secondary schools: educational laws from 1836 to 1933. BJHS Themes, 3:147–165, 2018. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/observation-
versus-experimentation-in-
naturalhistory-teaching-in-
portuguese-secondary-
schools-educational-laws-
from-1836-to-1933/
C4B7D85E400F237BFBCAA90BED3AA7A5.

Grote:2020:TKE

[Gro20] Mathias Grote. Total knowledge? Encyclopedic handbooks in the twentieth-century chemical and life sciences. BJHS Themes, 5::187–203, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/total-knowledge-

                                        encyclopedic-handbooks-in-
                                        the-twentiethcentury-
                                        chemical-and-life-sciences/
                                        8A4BB49EC370B5351F35A941C8DCE84A.

Hale:2021:CDS

[Hal21] Piers J. Hale. Charles Darwin, sexual selection and the evolution of other-regarding ethics. BJHS Themes, 6:157–177, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/charles-
                                        darwin-sexual-selection-
                                        and-the-evolution-of-
                                        otherregarding-ethics/
                                        9914E7B322270C98C1F4D2635904FAB2.

Hamlin:2021:DBP

[Ham21] Kimberly A. Hamlin. Darwin’s bawdy: the popular, gendered and radical reception of the Descent of Man in the US, 1871– 1910. BJHS Themes, 6:115–131, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/darwins-bawdy-the-
                                        popular-gendered-and-
                                        radical-reception-of-the-
                                        descent-of-man-in-the-us-
                                        18711910/
                                        53E66A9732DABB55D94DF5AB6A339855.

Hanson:2020:UEP

[Han20] Marta Hanson. From under the elbow to pointing to the palm: Chinese metaphors for learning medicine by the book (fourth– fourteenth centuries). BJHS Themes, 5::75–92, 2020. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/from-under-
the-elbow-to-pointing-to-
the-palm-chinese-metaphors-
for-learning-medicine-by-
the-book-fourthfourteenth-
centuries/
FAA61E6AF2FBCBD6758998A1BCB1B633.

Holloway:2017:BHB

[HB17] Lewis Holloway and Christopher Bear. Bovine and human becomings in histories of dairy technologies: robotic milking systems and remaking animal and human subjectivity. BJHS Themes, 2:215– 234, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https://www.

cambridge.org/core/product/
8336DA9ACBFA2A212D589B675605A2E6.

Hopwood:2019:TEF

[Hop19] Nick Hopwood. The tragedy of the emeritus and the fates of anatomical collections: Alfred Benninghoff’s memoir of Ferdinand Count Spee. BJHS Themes, 4:169–194, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/tragedy-of-
the-emeritus-and-the-fates-
of-anatomical-collections-
alfred-benninghoffs-memoir-
of-ferdinand-count-spee/
92C595FBAB75D7AB3FC9F30005B502EB.

Jardine:2019:MLH

[Jar19] Boris Jardine. The museum in the lab: historical practice in the experimental sciences at Cambridge, 1874–1936. BJHS Themes, 4:245–271, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/museum-in-
                                        the-lab-historical-
                                        practice-in-the-
                                        experimental-sciences-at-
                                        cambridge-18741936/
                                        C7DC52E93C3B9A6661F836BF3C1E2F73.

Jardine:2020:BIC

[Jar20] Boris Jardine. The book as instrument: craft and technique in early modern practical mathematics. BJHS Themes, 5:: 111–129, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/book-as-instrument-
                                        craft-and-technique-in-
                                        early-modern-practical-
                                        mathematics/
                                        5B7689CDFCCFCE5EA006206C0993ED5C.

Jiang:2021:LAD

[Jia21] Lijing Jiang. The late ascent of Darwin’s Descent: exploring human evolution and women’s role for a new China, 1927–1965. BJHS Themes, 6:201–220, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-

themes/article/late-ascent-
of-darwins-descent-
exploring-human-evolution-
and-womens-role-for-a-new-
china-19271965/
01BA8961286F3D6F3A4EBAA4D310ECB6.

Jardine:2019:HCE

[JKB19] Boris Jardine, Emma Kowal, and Jenny Bangham. How collections end: objects, meaning and loss in laboratories and museums. BJHS Themes, 4:1–27, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/how-
collections-end-objects-
meaning-and-loss-in-
laboratories-and-museums/
8BB44CBED3B7042D4481B77AA5613FA1.

Jones:2017:RIF

[Jon17] Karen R. Jones. Restor(y)ing the ‘fierce green fire’: animal agency, wolf conservation and environmental memory in Yellowstone National Park. BJHS Themes, 2:151–168, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
B296D4C4896DA91F40ED0479B81E6E66.

Kakaliouras:2019:RPK

[Kak19] Ann M. Kakaliouras. The repatriation of the Palaeoamericans: Kennewick Man/the Ancient One and the end of a non-Indian ancient North America. BJHS Themes, 4:79–98, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/
                                        repatriation-of-the-
                                        palaeoamericans-kennewick-
                                        manthe-ancient-one-and-the-
                                        end-of-a-nonindian-ancient-
                                        north-america/
                                        212A77EA4B4424BBBF898B9CB376C8E0.

Kaplan:2021:UBM

[Kap21] Judith R. H. Kaplan. ‘Unravelling Babel’: Mary LeCron Foster on the origins of language. BJHS Themes, 6:97–113, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/unravelling-
                                        babel-mary-lecron-foster-
                                        on-the-origins-of-language/
                                        C08C1F1F7EF2A1DB2A6183BDA223B879.

Kowal:2019:SDD

[Kow19] Emma Kowal. Spencer’s double: the decolonial afterlife of a postcolonial museum prop. BJHS Themes, 4:55–77, 2019. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/spencers-
                                        double-the-decolonial-
                                        afterlife-of-a-
                                        postcolonial-museum-prop/
                                        C4230DA0AF1B23CE04D15A7788E7727F.

Langlitz:2017:SPW

[Lan17] Nicolas Langlitz. Synthetic primatology: what humans and chimpanzees do in a Japanese laboratory and the African field. BJHS Themes, 2:101–125, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
5457633C714363D98577F2ED89281C12.

Leong:2020:LMB

[Leo20] Elaine Leong. Learning medicine by the book: reading and writing surgical manuals in early modern London. BJHS Themes, 5:: 93–110, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/learning-medicine-
by-the-book-reading-and-
writing-surgical-manuals-
in-early-modern-london/
241A28F41B456979D33FBF58910A78F1.

LOPez:2019:TCC

[Lóp19] Ana María Gómez López. On taphonomy: collages and collections at the Geiseltalmuseum. BJHS Themes, 4:195–214, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/on-
taphonomy-collages-and-
collections-at-the-
geiseltalmuseum/
8F3E3BE0280FC53F5449969E83D0BE8F.

Lewis:2016:SSL

[LS16] Michael Lewis and E. Elena Songster. Studying the snow leopard: reconceptualizing conservation across the China–India border. BJHS Themes, 1:169–198, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        992F367FE853AA276EB4B229FE71A716.

Marcon:2020:BFT

[Mar20a] Federico Marcon. The ‘book’ as fieldwork: ‘textual institutions’ and nature knowledge in early modern Japan. BJHS Themes, 5::131–148, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/book-as-fieldwork-
                                        textual-institutions-and-
                                        nature-knowledge-in-early-
                                        modern-japan/
                                        537F9296139E71D87DD189258E7D2C10.

Martelli:2020:AHG

[Mar20b] Matteo Martelli. Ancient handbooks and Graeco–Egyptian collections of alchemical recipes. BJHS Themes, 5::39–55, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/ancient-
                                        handbooks-and-
                                        graecoegyptian-collections-
                                        of-alchemical-recipes/
                                        3F2A7926F9D784A43E5170E0BA50AFB5.

Meister:2020:ENL

[Mei20] Anna-Maria Meister. Ernst Neufert’s ‘Lebensgestaltungslehre’: formatting life beyond the built. BJHS Themes, 5::167–185, 2020. CODEN ????

ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/ernst-neuferts-
lebensgestaltungslehre-
formatting-life-beyond-the-
built/
6626A6B18EDEB920F223AD37C393D4A4.

Milam:2021:EDS

[Mil21] Erika Lorraine Milam. The evolution of Darwinian sexualities. BJHS Themes, 6:133–155, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/evolution-
of-darwinian-sexualities/
7754D924656733682771BAC1192B527E.

Milam:2021:DDR

[MS21] Erika Lorraine Milam and Suman Seth. Descent of Darwin: race, sex, and human nature. BJHS Themes, 6:1–8, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL

https://www.cambridge.org/
core/journals/bjhs-themes/
article/descent-of-darwin-
race-sex-and-human-nature/
9C3B471326CE177DA4D33974D59D1991.

Mukharji:2021:DBS

[Muk21] Projit Bihari Mukharji. Darwin’s bulbuls: South Asian cultures of bird fighting and Darwin’s theory of sexual selection. BJHS Themes, 6:63–79, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL

https://www.cambridge.org/

                                        core/journals/bjhs-themes/
                                        article/darwins-bulbuls-
                                        south-asian-cultures-of-
                                        bird-fighting-and-darwins-
                                        theory-of-sexual-selection/
                                        CC07C2B458A9C51524F2FCD4DD2CC075.

Muller-Wille:2020:PSH

[MWP20] Staffan Müller-Wille and Giuditta Parolini. Punnett squares and hybrid crosses: how Mendelians learned their trade by the book. BJHS Themes, 5::149–165, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/punnett-
                                        squares-and-hybrid-crosses-
                                        how-mendelians-learned-
                                        their-trade-by-the-book/
                                        18A1CE37A6EE536CC1CE1D4FF6FF3174.

Nelson:2017:WDD

[Nel17] Amy Nelson. What the dogs did: animal agency in the Soviet manned space flight programme. BJHS Themes, 2:79–99, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        12AACC1131EE4DDDDF7FDED0B68341E7.

Nielsen:2018:IPP

[Nie18] Kristian H. Nielsen. Ideas, politics and practices of integrated science teaching in the global Cold War. BJHS Themes, 3: 167–189, 2018. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

                                        //www.cambridge.org/core/

journals/bjhs-themes/
article/ideas-politics-and-
practices-of-integrated-
science-teaching-in-the-
global-cold-war/
8E3F78E182BE82F677411A943E3C90B8.

Pettit:2017:GCM

[Pet17] Michael Pettit. The great cat mutilation: sex, social movements and the utilitarian calculus in 1970s New York City. BJHS Themes, 2:57–78, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
9AD806C16AE7C6DE26F1E1D4B4E41C7B.

Philip:2016:SHP

[Phi16] Kavita Philip. Speculative histories: Photo essay. BJHS Themes, 1:239–248, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
E0050D0B1AB800B54B938D43A17F6734.

Phalkey:2016:SGC

[PL16] Jahnavi Phalkey and Tong Lam.

Science of giants: China and India in the twentieth century. BJHS Themes, 1:1–11, January 1, 2016. ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/article/science-
of-giants-china-and-india-
in-the-twentieth-century/
64C68A635054E4D6394C953A2978695C.

Porter:2019:CEC

[Por19] Dahlia Porter. Catalogues for an entropic collection: losses, gains and disciplinary exhaustion in the Hunterian Museum, Glasgow. BJHS Themes, 4:215–243, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/catalogues-for-an-
                                        entropic-collection-losses-
                                        gains-and-disciplinary-
                                        exhaustion-in-the-
                                        hunterian-museum-glasgow/
                                        B62A760128598CCA4BABC0FB9C34F39E.

Phalkey:2016:PST

[PW16] Jahnavi Phalkey and Zuoyue Wang. Planning for science and technology in China and India. BJHS Themes, 1:83–113, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        ADEE4E1AE59568A0B053E9E39AD7F6E2.

Radick:2017:AAA

[Rad17] Gregory Radick. Animal agency in the age of the Modern Synthesis: W. H. Thorpe’s example. BJHS Themes, 2:35–56, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/product/
                                        90193B4E6CEF7CD36CAE54F552C173BD.

Rampling:2020:RAG

[Ram20] Jennifer M. Rampling. Reading alchemically: guides to ‘philosophical’ practice in early modern England. BJHS Themes, 5::57–74, 2020. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/reading-
alchemically-guides-to-
philosophical-practice-in-
early-modern-england/
D4BB9ECC77207496D0051522F4553ACF.

Reardon:2019:CEE

[Rea19] Jenny Reardon. Commentary:

ends everlasting. BJHS Themes, 4:283–291, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/commentary-ends-
everlasting/
FD29C780BA69D75B78ED9C2C748365B6.

Rees:2017:AAH

[Ree17a] Amanda Rees. Animal agents?

Historiography, theory and the history of science in the Anthropocene. BJHS Themes, 2:1–10, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

tronic).  URL https://www.
cambridge.org/core/product/
CF8362D3CAA3B027AC275C169EB76366.

Rees:2017:WAP

[Ree17b] Amanda Rees. Wildlife agencies: practice, intentionality and history in twentieth-century animal field studies. BJHS Themes, 2:127–149, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        78DEE9607A5E062959337581063AA128.

Rocha:2016:HDL

[Roc16] Leon Antonio Rocha. How deep is love? The engagement with India in Joseph Needham’s historiography of China. BJHS Themes, 1:13–41, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        3E411804F1E3DE83B951A3FF0D4673EF.

Roque:2019:BRC

[Roq19] Ricardo Roque. The blood that remains: card collections from the colonial anthropological missions. BJHS Themes, 4:29–53, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

                                        tronic).  URL https://www.
                                        cambridge.org/core/
                                        journals/bjhs-themes/
                                        article/blood-that-remains-
                                        card-collections-from-the-
                                        colonial-anthropological-
                                        missions/
                                        0E3AA84629B838FE911ACDC8635DEFF4.

Roy:2018:SCC

[Roy18]  Gautam Chando Roy.  Sci-

ence for children in a colonial context: Bengali juvenile magazines, 1883–1923. BJHS Themes, 3:43–72, 2018. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/science-for-
                                        children-in-a-colonial-

context-bengali-juvenile-
magazines-18831923/
9E96B955D9A48C2DB386CE2873BD5F86.

Radick:2021:LMC

[RS21] Gregory Radick and Mark Steadman. Of lice and men: Charles Darwin, Henry Denny and the evidence for the human races as varieties or species. BJHS Themes, 6:81–95, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/of-lice-and-
men-charles-darwin-henry-
denny-and-the-evidence-for-
the-human-races-as-
varieties-or-species/
CD63D99818210975B0827E5CB27E74A2.

Saha:2017:CEA

[Sah17] Jonathan Saha. Colonizing elephants: animal agency, undead capital and imperial science in British Burma. BJHS Themes, 2:169–189, 2017. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
34A559317F45848054C76C384BEEB8F3.

Sayer:2017:MMR

[Say17] Karen Sayer. The ‘modern’ management of rats: British agricultural science in farm and field during the twentieth century. BJHS Themes, 2:235–263, 2017. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

tronic).  URL https://www.

                                        cambridge.org/core/product/
                                        D82957541494D97C480EDF95859C0F15.

Seth:2021:CSD

[Set21] Suman Seth. ‘Constitutions selection’: Darwin, race and medicine. BJHS Themes, 6:25–43, 2021. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/
                                        constitutions-selection-
                                        darwin-race-and-medicine/
                                        FA3E4ADA514483DD89DC8EDAFBE93C15.

Sheldon:2021:SSR

[She21] Myrna Perez Sheldon. Sexual selection as race making. BJHS Themes, 6:9–23, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/sexual-
                                        selection-as-race-making/
                                        5FFC793521F988A4900E5B61F8F9F9DB.

Siddiqi:2016:AGH

[Sid16] Asif Siddiqi. Another global history of science: making space for India and China. BJHS Themes, 1:115–143, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/product/
                                        AFA74873CCA3A5927CF5E4BA78451DD6.

Somerset:2018:PPP

[Som18] Richard Somerset. Palaeontological pedagogues of the 1830s: the prehistory of the ‘history of life’ genre. BJHS Themes, 3:15–42, 2018. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/palaeontological-
pedagogues-of-the-1830s-
the-prehistory-of-the-
history-of-life-genre/
16A186FF5AB9B20EE99012D59D8135A0.

Sommer:2021:MAP

[Som21] Marianne Sommer. The meaning of absence: the primate tree that did not make it into Darwin’s The Descent of Man. BJHS Themes, 6:45–61, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/journals/bjhs-
themes/article/meaning-of-
absence-the-primate-tree-
that-did-not-make-it-into-
darwins-the-descent-of-man/
C294250D24E41516873CBB87CA83DA2F.

Saha:2016:GRE

[SS16] Madhumita Saha and Sigrid Schmalzer. Green-revolution epistemologies in China and India: technocracy and revolution in the production of scientific knowledge and peasant identity. BJHS Themes, 1:145–167, 2016. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

URL https://www.cambridge.
org/core/product/
38253E1DA101DEBAA1D0B65A9193E67E.

Skinner:2019:WED

[SW19] David Skinner and Matthias Wienroth. Was this an ending?

The destruction of samples and deletion of records from the UK Police National DNA Database. BJHS Themes, 4:99–121, 2019. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/was-this-an-
                                        ending-the-destruction-of-
                                        samples-and-deletion-of-
                                        records-from-the-uk-police-
                                        national-dna-database/
                                        B9454A08928AAE907FB0C8FF7103CFA3.

Wu:2016:INW

[Wu16] Pin-Hsien Wu. Investigating nature within different discursive and ideological contexts: case studies of Chinese and Indian coal capitals. BJHS Themes, 1:199– 220, 2016. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https://www.

                                        cambridge.org/core/product/
                                        3844E87C0B418000AD4714FFE6F70A90.

Zakariya:2021:EAP

[Zak21] Nasser Zakariya. Evolutionary antagonisms and the progress of three categories of traits. BJHS Themes, 6:179–200, 2021. CO-DEN ???? ISSN 2058-850X (print), 2056-354X (electronic).

                                        URL https://www.cambridge.
                                        org/core/journals/bjhs-
                                        themes/article/
                                        evolutionary-antagonisms-
                                        and-the-progress-of-three-
                                        categories-of-traits/
                                        1433754AE1AEB578551710F98219160A.

Zilhao:2018:TSC

[Zil18] Isabel Zilhão. Taking science to the countryside: fictionalizing the country through novels for young people in early twentiethcentury Portugal. BJHS Themes, 3:73–103, 2018. CODEN ???? ISSN 2058-850X (print), 2056-354X (electronic). URL https:

//www.cambridge.org/core/
journals/bjhs-themes/
article/taking-science-to-
the-countryside-
fictionalizing-the-country-
through-novels-for-young-
people-in-early-
twentiethcentury-portugal/
FA364C9D94FE200597BEFEECF20CCFB6.

Zilhao:2018:WSC

[ZN18] Isabel Zilhão and Kristian H.

Nielsen. Worlds of science for children and young people, 1830– 1991. BJHS Themes, 3:1–14, 2018. CODEN ???? ISSN 2058-850X (print), 2056-354X (elec-

tronic).  URL https://www.
cambridge.org/core/
journals/bjhs-themes/
article/worlds-of-science-
for-children-and-young-
people-18301991/
4B1B1414646DC568D54179828B0B3462.